OpenStudy (anonymous):

What are the solutions of : e. 2(t -1) + 3t < 2 f. 4n +2 < 7(n – 13) g. 14 – 6n ≤ 2 (7-3n) h. 4p + 3 > 9p +6 -5p

OpenStudy (texaschic101):

(e) 2(t - 1) + 3t < 2 2t - 2 + 3t < 2 5t - 2 < 2 5t < 2 + 2 5t < 4 t < 4/5 (f) 4n + 2 < 7(n - 13) 4n + 2 < 7n - 91 4n - 7n < -91 - 2 - 3n < - 93 -3n/-3 < -93/-3 n > 31 (g) 14 - 6n <= 2 (7 - 3n) 14 - 6n <= 14 - 6n -6n + 6n <= 14 - 14 0 = 0 infinite solutions (h) 4p + 3 > 9p + 6 - 5p 4p + 3 > 4p + 6 4p - 4p > 6 - 3 0 > 3 (not true) no solution